Shadee Corp. expects to sell 560 sun visors in May and 330 in June. Each visor sells for $11. Shadee’s beginning and ending finished goods inventories for May are 60 and 45 units, respectively. Ending finished goods inventory for June will be 60 units.

It expects the following unit sales for the third quarter:      

July 530
August 470
September 460


Sixty percent of Shadee’s sales are cash. Of the credit sales, 52 percent is collected in the month of the sale, 35 percent is collected during the following month, and 13 percent is never collected.

Required:

Calculate Shadee’s total cash receipts for August and September.
Answer #1

Calculate total cash receipt

August September
Cash sales 470*11*60% = 3102 460*11*60% = 3036
Collection from credit sales
July sales 530*11*40%*35% = 816.20
August sales 470*11*40%*52% = 1075.36 470*11*40%*35% = 723.80
September sales 460*11*40%*52% = 1052.48
Total cash receipts 4993.56 4812.28